OpenStudy (anonymous):

Which of the following is a solution to log 4 x + log 4 (x-3) = 1? A) x=6/5 B) x=2 C) x=-3 D) x=4

OpenStudy (cwrw238):

log4 x(x-3) = 1 x(x-3) = 4^1 x^2 - 3x - 4 = 0 x = 4 or- 1 x = 4 is correct